At the instant the traffic light turns green, a car that has been waiting at an intersection starts ahead with a constant acceleration of 3.2m/s2. At the same instant a truck, traveling with a constant speed of 20m/s overtakes and passes the car. How far beyond its starting point does the car overtake the truck in meters?
1
Expert's answer
2020-07-28T10:27:20-0400

"S_1=S_2"


"vt=at^2/2\to t=2v/a=2\cdot20/3.2=12.5(s)"


"S=at^2/2=3.2\cdot12.5^2/2=250(m)". Answer